You are given a 12 g sample of a really smelly compound and told by your boss that you need to perform combustion analysis to determine the empirical formula. He tells you that the compound doesn’t contain oxygen. Holding your breath, you run the combustion analysis and find that you produced 21.41 g CO2, 14.59 g H2O and 17.51 g of N2O5.

What is the empirical formula of the original compound?
Answer #1

Answer :-

Given:-

mass of CO2 = 21.41 g

mass of H2O = 14.59 g

mass of N2O5 = 17.51 g

wt. of sample = 12 g

As we know that

molar mass of CO2 = molar mass of C + 2 molar mass of O

molar mass of CO2 = 12 + 2 16

molar mass of CO2 = 12 + 32

molar mass of CO2 = 44 g /mol

similarly

molar mass of N2O5   = 2 molar mass of N + 5 molar mass of O

molar mass of N2O5   = 2 14 + 5 16

molar mass of N2O5   = 28 + 80

molar mass of N2O5   = 108 g /mol

similarly

molar mass of H2O  = 2 molar mass of H + molar mass of O

molar mass of H2O  = 2 + 16

molar mass of H2O  = 18 g / mol

Also we know that

44 g CO2 contained = 12 g C

1 g CO2 contained = 12 / 44 g C

then

21.41 g CO2 contained = 12 21.41 / 44 g C

21.41 g CO2 contained = 256.92 / 44 g C

21.41 g CO2 contained = 5.84 g C

Also we know that

% of C = wt. of C 100 / wt. of sample

% of C = 5.84 g 100 / 12 g

% of C = 584 / 12

% of C = 48.67 %

similarly

18 g H2O contained = 2 g H

1 g H2O contained = 2 / 18 g H

then

14.59 g H2O contained = 2 14.59 / 18 g H

14.59 g H2O contained = 29.18 / 18 g H

14.59 g H2O contained = 1.62 g H

Also we know that

% of H = wt. of H 100 / wt. of sample

% of H = 1.62 g 100 / 12 g

% of H = 162 / 12

% of H = 13.5 %

similarly

108 g N2O5 contained = 28 g N

1 g  N2O5 contained = 28 / 108 g N

then

17.51 g N2O5 contained = 28 17.51/ 108 g N

17.51 g N2O5 contained = 490.28 / 108 g N

17.51 g N2O5 contained = 4.54 g N

Also we know that

% of N = wt. of N 100 / wt. of sample

% of N = 4.54 g 100 / 12 g

% of N = 454 / 12

% of N = 37.83 %

therefore

Elements % Atomic wt. Atomic ratio HCF Atoms present
C 48.67 12 48.67/12 = 4.06 4.06 / 2.70 = 1.5   1.5 2 = 3
H 13.5 1 13.5 / 1 = 13.50 13.50 / 2.70 = 5   5 2 = 10
N 37.83 14 37.83 /14 = 2.70   2.70 / 2.70 = 1 1 ​​​​​​​ 2 = 2

So the empirical formula of the original compound is C3H10N2 (i.e the answer)